# PPA-Game: Characterizing and Learning Competitive Dynamics Among Online Content Creators

### v1

@@ A.2 High Probability of PNE Existence
For every randomly generated game configuration, considering the computational complexity of evaluating all strategy profiles, we trace an improvement path starting from the strategy profile $\bm{\pi}(0)=(0,0,\dots,0)$. If this improvement path concludes within finite steps, it implies the existence of a PNE. Out of the $5,000,000$ randomly generated game configurations in each distribution (i.e., totaling $20,000,000$ randomly generated configurations), only one exhibited an infinite improvement path. This suggests the existence of a PNE in the remaining $19,999,999$ configurations. A comprehensive evaluation of strategy profiles confirmed the non-existence of a PNE in the only exceptional case, concisely illustrated in Appendix A.1.
@@ Appendix C Omitted Proofs
Since $\bm{\pi}(s+1)$ is an improvement step of $\bm{\pi}(s)$, player $j$ earns a larger payoff by deviation in this step, leading to the fact that $w_{j,k}>0$ and the right-hand side of Equation (18) is well-defined. Furthermore, by the definition of $Q_{k}(\ell)$, we have
$W_{k}(\bm{\pi}(s+1))=W_{k}(\bm{\pi}(s))+w_{j,k}\geq Q_{k}(\ell)+w_{j,k}.$ (19)
@@
$\mathcal{U}_{j}(\bm{\pi}(s+1))=\frac{\mu_{k}w_{j,k}}{W_{k}(\bm{\pi}(s+1))}\leq\frac{\mu_{k}w_{j,k}}{Q_{k}(\ell)+w_{j,k}}.$ (20)
@@
[…]
@@
$\pi_{j(s)}(s)=k\quad\text{and}\quad Q_{k}(\ell)=W_{k}(\bm{\pi}(s+1)).$ (22)
@@
$\mathcal{U}_{j}(\bm{\pi}(s))=\frac{\mu_{k}w_{j,k}}{W_{k}(\bm{\pi}(s))}=\frac{\mu_{k}w_{j,k}}{W_{k}(\bm{\pi}(s+1))+w_{j,k}}=\frac{\mu_{k}w_{j,k}}{Q_{k}(\ell)+w_{j,k}}.$ (23)
@@
$\frac{\mu_{k}}{Q_{k}(\ell)+\epsilon_{0}}<\frac{\mu_{k^{\prime}}}{Q_{k^{\prime}}(\ell)+\epsilon_{0}}.$ (24)
@@
$\mathcal{U}_{j}(\bm{\pi}(s^{\prime}))=\frac{\mu_{k}w_{j,k}}{Q_{k}(\ell)+w_{j,k}}.$ (25)
@@
$\frac{\mu_{k}w_{j,k}}{Q_{k}(\ell)+w_{j,k}}=\mathcal{U}_{j}(\bm{\pi}(s^{\prime}))<\mathcal{U}_{j}(\bm{\pi}(s^{\prime}+1))\leq\frac{\mu_{k^{\prime}}w_{j,{k^{\prime}}}}{Q_{k^{\prime}}(\ell)+w_{j,k^{\prime}}},$ (26)
@@
First, we claim that $\mu_{k}\leq\mu_{k^{\prime}}$ by contradiction. Otherwise, we assume $\mu_{k}>\mu_{k^{\prime}}$ and we obtain that $\mu_{k}/\mu_{k^{\prime}}>N_{0}/\epsilon_{0}$ by the assumption. Based on Equation (26), we have:
$\frac{\mu_{k^{\prime}}}{\mu_{k}}>\frac{w_{j,k}}{Q_{k}(\ell)+w_{j,k}}\left/\frac{w_{j,k^{\prime}}}{Q_{k^{\prime}}(\ell)+w_{j,k^{\prime}}}\right..$ (27)
@@
$\frac{w_{j,k}}{Q_{k}(\ell)+w_{j,k}}\geq\frac{\epsilon_{0}}{N_{0}}\quad\text{and}\quad\frac{w_{j,k^{\prime}}}{Q_{k^{\prime}}(\ell)+w_{j,k^{\prime}}}\leq 1.$ (28)
@@
$\frac{\mu_{k^{\prime}}}{\mu_{k}}>\frac{w_{j,k}}{Q_{k}(\ell)+w_{j,k}}\left/\frac{w_{j,k^{\prime}}}{Q_{k^{\prime}}(\ell)+w_{j,k^{\prime}}}\right.\geq\frac{\epsilon_{0}}{N_{0}}>\frac{\mu_{k^{\prime}}}{\mu_{k}}.$ (29)
@@
$\frac{\mu_{k^{\prime}}}{\mu_{k}}>\frac{N_{0}}{\epsilon_{0}}\geq\frac{W_{k^{\prime}}(\bm{\pi}(s^{\prime}+1))}{Q_{k}(\ell)+\epsilon_{0}}\geq\frac{Q_{k^{\prime}}(\ell)+\epsilon_{0}}{Q_{k}(\ell)+\epsilon_{0}}.$ (30)
@@
$\ell=\left(\bm{\pi}(1),\bm{\pi}(2),\dots,\bm{\pi}(t),\bm{\pi}(t+1)=\bm{\pi}(1)\right).$ (31)
@@
$\frac{\mu_{k_{1}}}{Q_{k_{1}}(\ell)+\epsilon_{0}}<\frac{\mu_{k_{2}}}{Q_{k_{2}}(\ell)+\epsilon_{0}}<\dots<\frac{\mu_{k_{t+1}}}{Q_{k_{t+1}}(\ell)+\epsilon_{0}}.$ (32)
Since $k_{i}$ ($11\leq i\leq t+$ 1) can only take value in $\{1,2,\dots,t\}$, there must exists two indices $i_{1}\neq i_{2}$ such that $k_{i_{1}}=k_{i_{2}}$. This leads to a contradiction since Equation (32) suggests that $\forall 1\leq i_{1},i_{2}\leq t+1$, $k_{i_{1}}\neq k_{i_{2}}$. Now we can conclude that all improvement paths end in finite steps and the PNE exists. ∎
@@
$\left(\bm{\pi}(1),\bm{\pi}(2),\dots,\bm{\pi}(t),\bm{\pi}(t+1)=\bm{\pi}(1)\right).$ (34)
@@
$\Phi(\bm{\pi}(s))=\mathrm{Sort}\left(\frac{\mu_{k}}{W_{k}(\bm{\pi}(s))}\right)_{k=1}^{K}.$ (35)
@@
$\frac{\mu_{\pi_{j}(s)}}{W_{\pi_{j}(s)}(\bm{\pi}(s+1))}\geq\frac{\mu_{\pi_{j}(s)}}{W_{\pi_{j}(s)}(\bm{\pi}(s))}\quad\text{and}\quad\frac{\mu_{\pi_{j}(s+1)}}{W_{\pi_{j}(s+1)}(\bm{\pi}(s))}>\frac{\mu_{\pi_{j}(s+1)}}{W_{\pi_{j}(s+1)}(\bm{\pi}(s+1))}.$ (37)
Since $\Phi(\bm{\pi}(s+1))$ differs from $\Phi(\bm{\pi}(s))$ only on the items w.r.t. $k=\pi_{j}(s)$ and $k=\pi_{j}(s+1)$, by examining Equation (36) and Equation (37), it is easy to check that $\Phi(\bm{\pi}(s+1))$ is greater than $\Phi(\bm{\pi}(s))$ by lexicographic order. As a result,
$\Phi(\bm{\pi}(1))=\Phi(\bm{\pi}(t+1))\succ\Phi(\bm{\pi}(t))\succ\dots\succ\Phi(\bm{\pi}(1)),$ (38)
@@
$\left(\bm{\pi}(1),\bm{\pi}(2),\dots,\bm{\pi}(t),\bm{\pi}(t+1)=\bm{\pi}(1)\right).$ (39)
@@
$\frac{w_{j,\pi_{j}(s+1)}}{W_{\pi_{j}(s+1)}(\bm{\pi}(s+1))}=\frac{w_{j,k(s)}}{w_{j,k(s)}+W_{k(s)}\left(\bm{\pi}(s)\right)}>\frac{w_{j,\pi_{j}(s)}}{W_{\pi_{j}(s)}(\bm{\pi}(s))}.$ (40)
@@
$-\frac{W_{\pi_{j}(s+1)}\left(\bm{\pi}(s)\right)}{W_{\pi_{j}(s+1)}(\bm{\pi}(s+1))}=-\frac{W_{k(s)}\left(\bm{\pi}(s)\right)}{w_{j,k(s)}+W_{k(s)}\left(\bm{\pi}(s)\right)}>-\frac{W_{\pi_{j}(s)}(\bm{\pi}(s))-w_{j,\pi_{j}(s)}}{W_{\pi_{j}(s)}(\bm{\pi}(s))}=-\frac{W_{\pi_{j}(s)}(\bm{\pi}(s+1))}{W_{\pi_{j}(s)}(\bm{\pi}(s))}.$ (41)
@@
$\frac{W_{\pi_{j}(s)}(\bm{\pi}(s+1))}{W_{\pi_{j}(s)}(\bm{\pi}(s))}>\frac{W_{\pi_{j}(s+1)}\left(\bm{\pi}(s)\right)}{W_{\pi_{j}(s+1)}(\bm{\pi}(s+1))}.$ (42)
@@
$\Phi(\bm{\pi}(s))=\prod_{k\in\mathcal{K}}\left(\mathbb{I}[W_{k}(\bm{\pi}(s))=0]c_{0}+\mathbb{I}[W_{k}(\bm{\pi}(s))>0]W_{k}(\bm{\pi}(s))\right).$ (44)
By definition, we have $\Phi(s)>0$ for any $1\leq s\leq t$. Consider the change of the potential function for any round $s$ and $s+1$. Since $W_{k}(\bm{\pi}(s))$ differs from $W_{k}(\bm{\pi}(s+1))$ only when $k=\pi_{j(s)}(s)$ and $k=\pi_{j(s)}(s+1)$. Let $j=j(s)$. By Equation (42), we have $W_{\pi_{j}(s)}(\bm{\pi}(s)),W_{\pi_{j}(s+1)}(\bm{\pi}(s+1)),W_{\pi_{j}(s)}(\bm{\pi}(s+1))>0$. Consider the two cases when $W_{\pi_{j}(s+1)}\left(\bm{\pi}(s)\right)=0$ or $W_{\pi_{j}(s+1)}\left(\bm{\pi}(s)\right)>0$.
@@
    $\frac{\Phi(\bm{\pi}(s+1))}{\Phi(\bm{\pi}(s))}=\frac{W_{\pi_{j}(s)}(\bm{\pi}(s+1))W_{\pi_{j}(s+1)}(\bm{\pi}(s+1))}{W_{\pi_{j}(s)}(\bm{\pi}(s))\cdot c_{0}}>2$ (45)
@@
    $\frac{\Phi(\bm{\pi}(s+1))}{\Phi(\bm{\pi}(s))}=\frac{W_{\pi_{j}(s)}(\bm{\pi}(s+1))W_{\pi_{j}(s+1)}(\bm{\pi}(s+1))}{W_{\pi_{j}(s)}(\bm{\pi}(s))W_{\pi_{j}(s+1)}\left(\bm{\pi}(s)\right)}>1$ (46)
    by Equation (42).
@@
[…]
@@
Suppose the condition in Equation (48) does not hold for a Nash equilibrium $\bm{\pi}$. Then there must exist two indices, $k_{1}<k_{2}$ such that arm $k_{1}$ is not chosen by any player and arm $k_{2}$ is chosen by at least one player. Now any player that pulls arm $k_{2}$ can increase his reward by deviating to arm $k_{1}$, which leads to a contradiction since $\bm{\pi}$ is a Nash equilibrium. ∎
@@
$\left|\mathcal{U}_{j}(\bm{\pi})-\mathcal{U}^{\prime}_{j}(\bm{\pi})\right|$ $=\left|\frac{\left(\hat{\mu}_{j,k}-\mu_{k}\right)w_{j,k}}{\sum_{j^{\prime}=1}^{N}\mathbb{I}[\pi_{j^{\prime}}=k]w_{j^{\prime},k}}+\gamma_{j,k}\right|\leq\left|\frac{\left(\hat{\mu}_{j,k}-\mu_{k}\right)w_{j,k}}{\sum_{j^{\prime}=1}^{N}\mathbb{I}[\pi_{j^{\prime}}=k]w_{j^{\prime},k}}\right|+\left|\gamma_{j,k}\right|$ (49)
@@
On the one hand, suppose $\bm{\pi}$ is a Nash equilibrium of $\mathcal{G}$. By the definition of $\delta$ and $\delta^{\text{NE}}$ in Equation (13), we have that
$\forall j\in[N],\pi^{\prime}\in[K]\text{ and }k\neq\pi_{j},\quad\mathcal{U}_{j}(\bm{\pi})-\mathcal{U}_{j}(\pi^{\prime},\bm{\pi}_{-j})\geq\delta^{\text{NE}}\geq\delta.$ (50)
@@
$\mathcal{U}^{\prime}_{j}(\bm{\pi})-\mathcal{U}^{\prime}_{j}(\pi^{\prime},\bm{\pi}_{-j})$ $=\mathcal{U}_{j}(\bm{\pi})-\mathcal{U}_{j}(\pi^{\prime},\bm{\pi}_{-j})+\left(\mathcal{U}^{\prime}_{j}(\bm{\pi})-\mathcal{U}_{j}(\bm{\pi})\right)+\left(\mathcal{U}_{j}(\pi^{\prime},\bm{\pi}_{-j})-\mathcal{U}_{j}^{\prime}(\pi^{\prime},\bm{\pi}_{-j})\right)$ (51)
@@
On the other hand, suppose $\bm{\pi}$ is a Nash equilibrium of $\mathcal{G}^{\prime}$. If $\bm{\pi}$ is not a Nash equilibrium of $\mathcal{G}$, by Equation (13), there must exist a strategy profile $\bm{\pi}$, a player $j$ and a deviation $\pi^{\prime}\neq\pi_{j}$ such that
$\mathcal{U}_{j}(\pi^{\prime},\bm{\pi}_{-j})\geq\mathcal{U}_{j}(\bm{\pi})+\delta^{\text{NoNE}}\geq\mathcal{U}_{j}(\bm{\pi})+\delta.$ (52)
@@
$\mathcal{U}^{\prime}_{j}(\pi^{\prime},\bm{\pi}_{-j})-\mathcal{U}^{\prime}_{j}(\bm{\pi})$ $=\mathcal{U}_{j}(\pi^{\prime},\bm{\pi}_{-j})-\mathcal{U}_{j}(\bm{\pi})+\left(\mathcal{U}_{j}(\bm{\pi})-\mathcal{U}^{\prime}_{j}(\bm{\pi})\right)+\left(\mathcal{U}_{j}^{\prime}(\pi^{\prime},\bm{\pi}_{-j})-\mathcal{U}_{j}(\pi^{\prime},\bm{\pi}_{-j})\right)$ (53)
@@
$\forall k^{\prime}>k,\forall j\in[N],\pi_{j}\neq k^{\prime}\quad\text{and}\quad\forall k^{\prime}\leq k,\exists j\in[N],\pi_{j}=k^{\prime}.$ (54)
@@
$W_{\mathcal{G}^{\prime}}(\bm{\pi})$ $=\sum_{j=1}^{N}\mathcal{U}^{\prime}_{j}(\bm{\pi})=\sum_{j=1}^{N}\left(\frac{\hat{\mu}_{j,\pi_{j}}w_{j,\pi_{j}}}{\sum_{j^{\prime}=1}^{N}\mathbb{I}[\pi_{j^{\prime}}=\pi_{j}]w_{j^{\prime},\pi_{j}}}+\gamma_{j,\pi_{j}}\right)$ (55)

### v2

@@ A.2 High Probability of PNE Existence
For every randomly generated game configuration, considering the computational complexity of evaluating all strategy profiles, we trace an improvement path starting from the strategy profile $\bm{\pi}(0)=(0,0,\dots,0)$. If this improvement path concludes within finite steps, it implies the existence of a PNE. Out of the $5,000,000$ randomly generated game configurations in each distribution (i.e., totaling $20,000,000$ randomly generated configurations), only one exhibited an infinite improvement path. This suggests the existence of a PNE in the remaining $19,999,999$ configurations. A comprehensive evaluation of strategy profiles confirmed the non-existence of a PNE in the only exceptional case, concisely illustrated in Section A.1.
@@ Appendix C Omitted Proofs
Since $\bm{\pi}(s+1)$ is an improvement step of $\bm{\pi}(s)$, player $j$ earns a larger payoff by deviation in this step, leading to the fact that $w_{j,k}>0$ and the right-hand side of Lemma C.1 is well-defined. Furthermore, by the definition of $Q_{k}(\ell)$, we have
$W_{k}(\bm{\pi}(s+1))=W_{k}(\bm{\pi}(s))+w_{j,k}\geq Q_{k}(\ell)+w_{j,k}.$
@@
$\mathcal{U}_{j}(\bm{\pi}(s+1))=\frac{\mu_{k}w_{j,k}}{W_{k}(\bm{\pi}(s+1))}\leq\frac{\mu_{k}w_{j,k}}{Q_{k}(\ell)+w_{j,k}}.$
@@
$\mathcal{U}_{j}\left(\bm{\pi}(s)\right)=\frac{\mu_{k}w_{j,k}}{Q_{k}(\ell)+w_{j,k}}\quad\text{where}\quad j=j(s).$
@@
$\pi_{j(s)}(s)=k\quad\text{and}\quad Q_{k}(\ell)=W_{k}(\bm{\pi}(s+1)).$
@@
$\mathcal{U}_{j}(\bm{\pi}(s))=\frac{\mu_{k}w_{j,k}}{W_{k}(\bm{\pi}(s))}=\frac{\mu_{k}w_{j,k}}{W_{k}(\bm{\pi}(s+1))+w_{j,k}}=\frac{\mu_{k}w_{j,k}}{Q_{k}(\ell)+w_{j,k}}.$
@@
$\frac{\mu_{k}}{Q_{k}(\ell)+\epsilon_{0}}<\frac{\mu_{k^{\prime}}}{Q_{k^{\prime}}(\ell)+\epsilon_{0}}.$
@@
$\mathcal{U}_{j}(\bm{\pi}(s^{\prime}))=\frac{\mu_{k}w_{j,k}}{Q_{k}(\ell)+w_{j,k}}.$
@@
$\frac{\mu_{k}w_{j,k}}{Q_{k}(\ell)+w_{j,k}}=\mathcal{U}_{j}(\bm{\pi}(s^{\prime}))<\mathcal{U}_{j}(\bm{\pi}(s^{\prime}+1))\leq\frac{\mu_{k^{\prime}}w_{j,{k^{\prime}}}}{Q_{k^{\prime}}(\ell)+w_{j,k^{\prime}}},$
@@
First, we claim that $\mu_{k}\leq\mu_{k^{\prime}}$ by contradiction. Otherwise, we assume $\mu_{k}>\mu_{k^{\prime}}$ and we obtain that $\mu_{k}/\mu_{k^{\prime}}>N_{0}/\epsilon_{0}$ by the assumption. Based on Section C.1, we have:
$\frac{\mu_{k^{\prime}}}{\mu_{k}}>\frac{w_{j,k}}{Q_{k}(\ell)+w_{j,k}}\left/\frac{w_{j,k^{\prime}}}{Q_{k^{\prime}}(\ell)+w_{j,k^{\prime}}}\right..$
@@
$\frac{w_{j,k}}{Q_{k}(\ell)+w_{j,k}}\geq\frac{\epsilon_{0}}{N_{0}}\quad\text{and}\quad\frac{w_{j,k^{\prime}}}{Q_{k^{\prime}}(\ell)+w_{j,k^{\prime}}}\leq 1.$
@@
$\frac{\mu_{k^{\prime}}}{\mu_{k}}>\frac{w_{j,k}}{Q_{k}(\ell)+w_{j,k}}\left/\frac{w_{j,k^{\prime}}}{Q_{k^{\prime}}(\ell)+w_{j,k^{\prime}}}\right.\geq\frac{\epsilon_{0}}{N_{0}}>\frac{\mu_{k^{\prime}}}{\mu_{k}}.$
@@
$\frac{\mu_{k^{\prime}}}{\mu_{k}}>\frac{N_{0}}{\epsilon_{0}}\geq\frac{W_{k^{\prime}}(\bm{\pi}(s^{\prime}+1))}{Q_{k}(\ell)+\epsilon_{0}}\geq\frac{Q_{k^{\prime}}(\ell)+\epsilon_{0}}{Q_{k}(\ell)+\epsilon_{0}}.$
@@
$\ell=\left(\bm{\pi}(1),\bm{\pi}(2),\dots,\bm{\pi}(t),\bm{\pi}(t+1)=\bm{\pi}(1)\right).$
@@
$\frac{\mu_{k_{1}}}{Q_{k_{1}}(\ell)+\epsilon_{0}}<\frac{\mu_{k_{2}}}{Q_{k_{2}}(\ell)+\epsilon_{0}}<\dots<\frac{\mu_{k_{t+1}}}{Q_{k_{t+1}}(\ell)+\epsilon_{0}}.$
Since $k_{i}$ ($11\leq i\leq t+$ 1) can only take value in $\{1,2,\dots,t\}$, there must exists two indices $i_{1}\neq i_{2}$ such that $k_{i_{1}}=k_{i_{2}}$. This leads to a contradiction since Section C.1 suggests that $\forall 1\leq i_{1},i_{2}\leq t+1$, $k_{i_{1}}\neq k_{i_{2}}$. Now we can conclude that all improvement paths end in finite steps and the PNE exists. ∎
@@
$\left(\bm{\pi}(1),\bm{\pi}(2),\dots,\bm{\pi}(t),\bm{\pi}(t+1)=\bm{\pi}(1)\right).$
@@
$\Phi(\bm{\pi}(s))=\mathrm{Sort}\left(\frac{\mu_{k}}{W_{k}(\bm{\pi}(s))}\right)_{k=1}^{K}.$
@@
$\frac{\mu_{\pi_{j}(s)}}{W_{\pi_{j}(s)}(\bm{\pi}(s+1))}\geq\frac{\mu_{\pi_{j}(s)}}{W_{\pi_{j}(s)}(\bm{\pi}(s))}\quad\text{and}\quad\frac{\mu_{\pi_{j}(s+1)}}{W_{\pi_{j}(s+1)}(\bm{\pi}(s))}>\frac{\mu_{\pi_{j}(s+1)}}{W_{\pi_{j}(s+1)}(\bm{\pi}(s+1))}.$
Since $\Phi(\bm{\pi}(s+1))$ differs from $\Phi(\bm{\pi}(s))$ only on the items w.r.t. $k=\pi_{j}(s)$ and $k=\pi_{j}(s+1)$, by examining Section C.2 and Section C.2, it is easy to check that $\Phi(\bm{\pi}(s+1))$ is greater than $\Phi(\bm{\pi}(s))$ by lexicographic order. As a result,
$\Phi(\bm{\pi}(1))=\Phi(\bm{\pi}(t+1))\succ\Phi(\bm{\pi}(t))\succ\dots\succ\Phi(\bm{\pi}(1)),$
@@
$\left(\bm{\pi}(1),\bm{\pi}(2),\dots,\bm{\pi}(t),\bm{\pi}(t+1)=\bm{\pi}(1)\right).$
@@
$\frac{w_{j,\pi_{j}(s+1)}}{W_{\pi_{j}(s+1)}(\bm{\pi}(s+1))}=\frac{w_{j,k(s)}}{w_{j,k(s)}+W_{k(s)}\left(\bm{\pi}(s)\right)}>\frac{w_{j,\pi_{j}(s)}}{W_{\pi_{j}(s)}(\bm{\pi}(s))}.$
@@
$-\frac{W_{\pi_{j}(s+1)}\left(\bm{\pi}(s)\right)}{W_{\pi_{j}(s+1)}(\bm{\pi}(s+1))}=-\frac{W_{k(s)}\left(\bm{\pi}(s)\right)}{w_{j,k(s)}+W_{k(s)}\left(\bm{\pi}(s)\right)}>-\frac{W_{\pi_{j}(s)}(\bm{\pi}(s))-w_{j,\pi_{j}(s)}}{W_{\pi_{j}(s)}(\bm{\pi}(s))}=-\frac{W_{\pi_{j}(s)}(\bm{\pi}(s+1))}{W_{\pi_{j}(s)}(\bm{\pi}(s))}.$
@@
$\frac{W_{\pi_{j}(s)}(\bm{\pi}(s+1))}{W_{\pi_{j}(s)}(\bm{\pi}(s))}>\frac{W_{\pi_{j}(s+1)}\left(\bm{\pi}(s)\right)}{W_{\pi_{j}(s+1)}(\bm{\pi}(s+1))}.$
@@
$\Phi(\bm{\pi}(s))=\prod_{k\in\mathcal{K}}\left(\mathbb{I}[W_{k}(\bm{\pi}(s))=0]c_{0}+\mathbb{I}[W_{k}(\bm{\pi}(s))>0]W_{k}(\bm{\pi}(s))\right).$
By definition, we have $\Phi(s)>0$ for any $1\leq s\leq t$. Consider the change of the potential function for any round $s$ and $s+1$. Since $W_{k}(\bm{\pi}(s))$ differs from $W_{k}(\bm{\pi}(s+1))$ only when $k=\pi_{j(s)}(s)$ and $k=\pi_{j(s)}(s+1)$. Let $j=j(s)$. By Section C.2, we have $W_{\pi_{j}(s)}(\bm{\pi}(s)),W_{\pi_{j}(s+1)}(\bm{\pi}(s+1)),W_{\pi_{j}(s)}(\bm{\pi}(s+1))>0$. Consider the two cases when $W_{\pi_{j}(s+1)}\left(\bm{\pi}(s)\right)=0$ or $W_{\pi_{j}(s+1)}\left(\bm{\pi}(s)\right)>0$.
@@
    $\frac{\Phi(\bm{\pi}(s+1))}{\Phi(\bm{\pi}(s))}=\frac{W_{\pi_{j}(s)}(\bm{\pi}(s+1))W_{\pi_{j}(s+1)}(\bm{\pi}(s+1))}{W_{\pi_{j}(s)}(\bm{\pi}(s))\cdot c_{0}}>2$
@@
    $\frac{\Phi(\bm{\pi}(s+1))}{\Phi(\bm{\pi}(s))}=\frac{W_{\pi_{j}(s)}(\bm{\pi}(s+1))W_{\pi_{j}(s+1)}(\bm{\pi}(s+1))}{W_{\pi_{j}(s)}(\bm{\pi}(s))W_{\pi_{j}(s+1)}\left(\bm{\pi}(s)\right)}>1$
    by Section C.2.
@@
Proof.
In the PPA-Game, the strategy profile that maximizes social welfare assigns all players to the top-$N$ resources if $N\leq K$, and if $N>K$, the top $K$ players select all $K$ resources. Consequently, the maximal social welfare is $\sum_{i=1}^{\min\{N,K\}}\mu_{i}$. Denote $W^{\max}=\sum_{i=1}^{\min\{N,K\}}\mu_{i}$ as the numerator of PoA given in Eq. 3.
Suppose $\bm{\pi}$ is a PNE. Let =$\mathcal{M}=\{1,2,\dots,\min\{N,K\}\}$ and $\mathcal{C}=\{j\in[K]:\exists i\in[N],\pi_{i}=j\}$. As a result, by the definition of social welfare and the utility function given in Eq. 1, we have that
$W^{\max}=\sum_{j\in\mathcal{M}}\mu_{j}\quad\text{and}\quad W_{\mathcal{G}}(\bm{\pi})=\sum_{j\in\mathcal{C}}\mu_{j}=\sum_{i=1}^{N}\mathcal{U}_{i}(\bm{\pi}).$
Note that since $\sum_{i=1}^{N}\mathcal{U}_{i}(\bm{\pi})=W_{\mathcal{G}}(\bm{\pi})$. There must exist one player $i^{*}$ such that $\mathcal{U}_{i}(\bm{\pi})\leq W_{\mathcal{G}}(\bm{\pi})/N$. Since $\bm{\pi}$ is a PNE, player $i^{*}$ will not benefit by deviation, which means that $\𝒞\forall j\in\mathcal{M}\backslash\mathcal{C}$, $\mu_{j}\leq W_{\mathcal{G}}(\bm{\pi})/N$. As a result,
$W^{\max}=\sum_{j\in\mathcal{M}}\mu_{j}\leq\sum_{j\in\mathcal{C}}\mu_{j}+\sum_{j\in\mathcal{M}\backslash\mathcal{C}}\mu_{j}\leq W_{\mathcal{G}}(\bm{\pi})+|\mathcal{M}\backslash\mathcal{C}|\cdot W_{\mathcal{G}}(\bm{\pi})/N.$
Note that we must have $1\in\mathcal{M}$ and $1\in\mathcal{C}$. Therefore, |\$|\mathcal{M}\backslash\mathcal{C}|\leq|\mathcal{M}|-1=\min\{N,K\}-1$. This gives us
$\frac{W^{\max}}{W_{\mathcal{G}}(\bm{\pi})}\leq 1+\frac{|\mathcal{M}\backslash\mathcal{C}|}{N}\leq 1+\frac{\min\{N,K\}-1}{N}.$
Now the claim follows. ∎
C.4 Proof of Theorem 4.1
@@
$\forall k^{\prime}>k,\forall j\in[N],\pi_{j}\neq k^{\prime}\quad\text{and}\quad\forall k^{\prime}\leq k,\exists j\in[N],\pi_{j}=k^{\prime}.$
@@
Suppose the condition in Lemma C.4 does not hold for a Nash equilibrium $\bm{\pi}$. Then there must exist two indices, $k_{1}<k_{2}$ such that arm $k_{1}$ is not chosen by any player and arm $k_{2}$ is chosen by at least one player. Now any player that pulls arm $k_{2}$ can increase his reward by deviating to arm $k_{1}$, which leads to a contradiction since $\bm{\pi}$ is a Nash equilibrium. ∎
@@
$\left|\mathcal{U}_{j}(\bm{\pi})-\mathcal{U}^{\prime}_{j}(\bm{\pi})\right|$ $=\left|\frac{\left(\hat{\mu}_{j,k}-\mu_{k}\right)w_{j,k}}{\sum_{j^{\prime}=1}^{N}\mathbb{I}[\pi_{j^{\prime}}=k]w_{j^{\prime},k}}+\gamma_{j,k}\right|\leq\left|\frac{\left(\hat{\mu}_{j,k}-\mu_{k}\right)w_{j,k}}{\sum_{j^{\prime}=1}^{N}\mathbb{I}[\pi_{j^{\prime}}=k]w_{j^{\prime},k}}\right|+\left|\gamma_{j,k}\right|$
@@
On the one hand, suppose $\bm{\pi}$ is a Nash equilibrium of $\mathcal{G}$. By the definition of $\delta$ and $\delta^{\text{NE}}$ in Eq. 8, we have that
$\forall j\in[N],\pi^{\prime}\in[K]\text{ and }k\neq\pi_{j},\quad\mathcal{U}_{j}(\bm{\pi})-\mathcal{U}_{j}(\pi^{\prime},\bm{\pi}_{-j})\geq\delta^{\text{NE}}\geq\delta.$
@@
$\mathcal{U}^{\prime}_{j}(\bm{\pi})-\mathcal{U}^{\prime}_{j}(\pi^{\prime},\bm{\pi}_{-j})$ $=\mathcal{U}_{j}(\bm{\pi})-\mathcal{U}_{j}(\pi^{\prime},\bm{\pi}_{-j})+\left(\mathcal{U}^{\prime}_{j}(\bm{\pi})-\mathcal{U}_{j}(\bm{\pi})\right)+\left(\mathcal{U}_{j}(\pi^{\prime},\bm{\pi}_{-j})-\mathcal{U}_{j}^{\prime}(\pi^{\prime},\bm{\pi}_{-j})\right)$
@@
On the other hand, suppose $\bm{\pi}$ is a Nash equilibrium of $\mathcal{G}^{\prime}$. If $\bm{\pi}$ is not a Nash equilibrium of $\mathcal{G}$, by Eq. 8, there must exist a strategy profile $\bm{\pi}$, a player $j$ and a deviation $\pi^{\prime}\neq\pi_{j}$ such that
$\mathcal{U}_{j}(\pi^{\prime},\bm{\pi}_{-j})\geq\mathcal{U}_{j}(\bm{\pi})+\delta^{\text{NoNE}}\geq\mathcal{U}_{j}(\bm{\pi})+\delta.$
@@
$\mathcal{U}^{\prime}_{j}(\pi^{\prime},\bm{\pi}_{-j})-\mathcal{U}^{\prime}_{j}(\bm{\pi})$ $=\mathcal{U}_{j}(\pi^{\prime},\bm{\pi}_{-j})-\mathcal{U}_{j}(\bm{\pi})+\left(\mathcal{U}_{j}(\bm{\pi})-\mathcal{U}^{\prime}_{j}(\bm{\pi})\right)+\left(\mathcal{U}_{j}^{\prime}(\pi^{\prime},\bm{\pi}_{-j})-\mathcal{U}_{j}(\pi^{\prime},\bm{\pi}_{-j})\right)$
@@
$\forall k^{\prime}>k,\forall j\in[N],\pi_{j}\neq k^{\prime}\quad\text{and}\quad\forall k^{\prime}\leq k,\exists j\in[N],\pi_{j}=k^{\prime}.$
@@
$W_{\mathcal{G}^{\prime}}(\bm{\pi})$ $=\sum_{j=1}^{N}\mathcal{U}^{\prime}_{j}(\bm{\pi})=\sum_{j=1}^{N}\left(\frac{\hat{\mu}_{j,\pi_{j}}w_{j,\pi_{j}}}{\sum_{j^{\prime}=1}^{N}\mathbb{I}[\pi_{j^{\prime}}=\pi_{j}]w_{j^{\prime},\pi_{j}}}+\gamma_{j,\pi_{j}}\right)$